OpenStudy (anonymous):

How to evaluate gamma function at n? \[\Gamma(n)=\int\limits_{0}^{\infty}e^{-t}t^{n-1}dt\] But how do we find gamma(1) or something (using this identity)?

OpenStudy (anonymous):

If you want gamma(1), you replace n with the value of 1. In general, you evaluate the integral (treating n as a constant and t as a variable). The result will be a function of n

OpenStudy (anonymous):

When the integral is evaluated, you will have gotten rid of the t's, so gamma(n) will have only n as a variable.

OpenStudy (anonymous):

Yes, but how do you evaluate the integral?

OpenStudy (anonymous):

Can't help you in that. I don't remember how you can solve it. Sorry

OpenStudy (dls):

Newton-Leibniz formula to the rescue !

OpenStudy (anonymous):

we ca forget the intergral and consider the defination \[\Gamma (n)=(n-1)!\] for any n we can evaluate this

OpenStudy (anonymous):

ok, but what about for say n = 1.5?

OpenStudy (anonymous):

how ever the improper integral also can be used to i will prove for \[\Gamma(1/2)\]

OpenStudy (anonymous):

\[\Gamma(\frac{ 1 }{ 2 })=\int\limits _0^{\infty}x^{1/2}e^{-x}dx\] \[\text{ let u=x^2 then } \implies 2 \int\limits_0^{\infty}e^{-u^2}du=2(\frac{ \sqrt{\pi} }{ 2 })=\sqrt{\pi}\]

OpenStudy (anonymous):

for fractions it gets weird but using the fact that e^-x^2 is sqrt of pi we can solve alot of fraction problems

OpenStudy (anonymous):

why let u = x^2?

OpenStudy (anonymous):

sry i meant x=u^2

OpenStudy (anonymous):

wich means u=x^1/2

OpenStudy (anonymous):

ok. How did you get from the integral to the next step?

OpenStudy (anonymous):

Before you answer that, does it involve anything beyond high school maths?

OpenStudy (anonymous):

that requires a new idea not offered in high school math ,its a special function critical to probability

OpenStudy (anonymous):

If it's not to difficult, would you care to expand?

OpenStudy (anonymous):

^too

OpenStudy (anonymous):

Oh, this is the error function...right, I've heard a bit about this.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when you substituted x = u^2, dx = 2u du right? So then where did the u go in the next step?

OpenStudy (anonymous):

you are correct \[2u du=dx\implies \] \[\int\limits _0^{\infty}e^{-x} \color{green}{x^{\frac{ -1 }{ 2 }}dx}=\int _0^{\infty} e^{-x}x^{-1/2}u2du=\int _0^{\infty} e^{-u^2}u^{-1}u2du\] another mistake we have \[x^{1-1/2}=x^{-1/2}\]\

OpenStudy (anonymous):

i had made a substitution mistake ealier its -1\2 not 1\2

OpenStudy (anonymous):

so th u cancel

OpenStudy (anonymous):

Yeah...I'm getting the gist of it. Can we generalise this to any value of n?

OpenStudy (anonymous):

i was wondering about that first we need to consider integers (already fine cos we will use factorial!)then we need to look at numers of the form p/q i guess with 1/2 we where just lucky in normal cases the x remaines as parrt of the intergral ,but we can always generalize for \[\Gamma(n/2)\]

OpenStudy (anonymous):

So you can do it? What if I gave you like n = 1913/17 or something? or what if n = pi?

OpenStudy (anonymous):

its particular values can be expressed in closed form for integer and half-integer arguments, but no simple expressions are known for the values at rational points in general

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Gamma_function

OpenStudy (anonymous):

So you can't do exact values for anything that isn't an integer or half integer?

OpenStudy (anonymous):

exaclty,as you say it ,mostly rational numbers are approaximated

OpenStudy (anonymous):

Interesting... Thanks heaps for all that by the way

OpenStudy (anonymous):

yw,are you really in high school

OpenStudy (anonymous):

yeah. Year 11

OpenStudy (anonymous):

great,good work and keep it up...sure you wont have problems when you meet up with integrals at college???

OpenStudy (anonymous):

Haha thanks! Actually I'm already doing integrals...it's part of the Australian curriculum